OpenStudy (lifeisadangerousgame):

@jdoe0001 Mind helping me again?

OpenStudy (lifeisadangerousgame):

I'm doing nearest integer this time, I know how to do it and everything ,I just don't understand the values part. Do I got one by one like this -1.5 = -2 -1.4 = -2 and so on or do I juts skip around how I please? -1.5 = -2 -1.2 = -4 -1 = -2 -0.5 = 0

OpenStudy (lifeisadangerousgame):

I'm really confused with all of this..if anyone could help me that'd be great..

OpenStudy (lifeisadangerousgame):

that's thep roblem btw

OpenStudy (lifeisadangerousgame):

well, since its nearest integer I know that it has to be rounded to the nearest even whole number, but I'm not sure if I have to do each number between -1.5 and 1.5 to get the graph I need

OpenStudy (lifeisadangerousgame):

Just -2 and 2? :/ will it look like this if I do that?